OpenStudy (anonymous):

Graph the data as a scatter plot. Draw a line through the “center-most” part of the data making sure to pass through at least two points of data. Use the two points on the line and the point-slope formula, y - y1 = m(x - x1), to write the equation of the line.

OpenStudy (anonymous):

HELPP

OpenStudy (anonymous):

Graph the data as a scatter plot. Draw a line through the “center-most” part of the data making sure to pass through at least two points of data. Use the two points on the line and the point-slope formula, y - y1 = m(x - x1), to write the equation of the line. Convert the equation of the line to slope-intercept form so that you can make predictions or answer questions about the data.

OpenStudy (anonymous):

@hkm91 Are you helping me or just staring at my question ?

OpenStudy (anonymous):

where the data is?

OpenStudy (anonymous):

you have to make it up Use your own numbers

OpenStudy (anonymous):

@hkm91 ^^

OpenStudy (anonymous):

post the data or should i take some random nos

OpenStudy (anonymous):

Random numbers @hkm91

OpenStudy (anonymous):

 data (1,1) (-1,-1) to draw a line

OpenStudy (anonymous):

and other no. near by this sort of method show data location and its density

OpenStudy (anonymous):

Omg thankyou o much

OpenStudy (anonymous):

But what are the numberrs ? @hkm91

OpenStudy (anonymous):

I mean where are they on the scatter plot ?

OpenStudy (anonymous):

@hkm91 ?

OpenStudy (anonymous):

@ThatHippieKid (1,1) (0,0) (-1,-1) (1,0)( 0,1) any decimals no. near by 1

OpenStudy (anonymous):

Alright thankyou, Do you mind drawing me another skatter plot so that i can copy it down ?

OpenStudy (anonymous):

Or could you do this Graph the data as a scatter plot. Draw a line through the “center-most” part of the data making sure to pass through at least two points of data. Use the two points on the line and the point-slope formula, y - y1 = m(x - x1), to write the equation of the line. Convert the equation of the line to slope-intercept form so that you can make predictions or answer questions about the data.

OpenStudy (anonymous):

Like i am not sure how to find out the point slope formula question or whatever, im confused. @hkm91

OpenStudy (anonymous):

do u know cartesion cordinates ploting?

OpenStudy (anonymous):

plot (1,1) (-1,-1) on the graph and join them this are in form of (x.y)

OpenStudy (anonymous):

No i dont know that, @hkm91

OpenStudy (anonymous):

@ThatHippieKid sorry than i cant help u

OpenStudy (anonymous):

Fine

OpenStudy (anonymous):

see this no. are written in the form of (x,y)

OpenStudy (anonymous):

I know that

OpenStudy (anonymous):

I know how to graph and everything i just cant figure out how to do this scatter plot. and it must be perfect because i need to bring my grade upin this class from a D to a B.

OpenStudy (anonymous):

ok then the in the data i have took two points are (1,1) and (-1,-1)

OpenStudy (anonymous):

equation of the line joining two points is given in the question

OpenStudy (anonymous):

where m is the slope.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

m =y2-y1/x2-x1

OpenStudy (anonymous):

I dont know what my graph is supposed to look like ?

OpenStudy (anonymous):

THANKSSS!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

How or what is the answer to that m = y2 + y1 and x2 - x1 ?

OpenStudy (anonymous):

@hkm91

OpenStudy (anonymous):

Can you draw me another map plot ? @hkm91

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) m =1

OpenStudy (anonymous):